OpenStudy (anonymous):

how to solve this quadratic eqution?find the range of values of "k" for which the roots of the equation x^2 + (2k+1)x+k^2 -3 = 0 , are not real. State the range of values of "k" for which the roots are real.

OpenStudy (anonymous):

Use the quadratic formula or complete the square. Then look at the term inside the radical. The radical of a real number is only real, if that number is not negative! So solve the inequality >= 0 for the term inside the radical.

OpenStudy (anonymous):

what is redicle?

OpenStudy (anonymous):

The square root!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The solutions of your quadratic equation are \[ -k-\frac{1}{2}\pm\frac{1}{2}\cdot\sqrt{4k+13} \].

OpenStudy (anonymous):

oh thankx i am tring to solve it :)

OpenStudy (anonymous):

k^2+(2 k+1) x+x^2-3 = 0 Expand out terms of the left hand side: k^2+2 k x+x^2+x-3 = 0 Subtract x^2+x-3 from both sides: k^2+2 k x = -x^2-x+3 Add x^2 to both sides: k^2+2 k x+x^2 = 3-x Factor the left hand side: (k+x)^2 = 3-x Take the square root of both sides: abs(k+x) = sqrt(3-x) Eliminate the absolute value: k+x = -sqrt(3-x) or k+x = sqrt(3-x) Subtract x from both sides: k = -sqrt(3-x)-x or k+x = sqrt(3-x) Subtract x from both sides: k = -sqrt(3-x)-x or k = sqrt(3-x)-x

OpenStudy (anonymous):

k=-+[sqrt(3-x)]-x k=-+[sqrt(3-x)] -x, for 0<x<3, k=Real and for x>3,k=imaginary ans